# Search for Dark Matter Annihilation to gamma-rays from SPT-SZ selected Galaxy Clusters

### v1

@@ Abstract
We search for dark matter annihilation from galaxy clusters in the energy range from 1-300 GeV using nearly 16 years of Fermi-LAT data. For this purpose, we use 350 galaxy clusters selected from the 2500 SPT-SZ survey. We model the dark matter distribution using the NFW profile for the main halo along with the Einasto profile for the substructure. The largest signal is seen for the cluster SPT-CL J2021-5257 with a significance of around . The best-fit dark matter mass and annihilation cross-section for this cluster are equal to GeV and for the annihilation channel. However, this central estimate is in conflict with the limits on annihilation cross-section from dwarf spheroidal galaxies and hence cannot be attributed to dark matter annihilation. Three other clusters show significance between , whereas all the remaining clusters show null results. The most stringent 95% c.l. upper limit, which we obtain among the clusters with significance is from SPT-CL J2300-5331, given by for a dark matter mass of 10 GeV corresponding to the annihilation channel.
@@ I Introduction
The dark matter problem is one of the most outstanding unsolved problems in Astrophysics Jungman et al. (1996); Bertone et al. (2005); Bertone and Hooper (2018); Bosma (2023). Although a large number of observations have pointed to a concordance model of the universe, which consists of about 25% cold dark matter Planck Collaboration et al. (2020), the identity of the cold dark matter candidate is unknown. One of the most well motivated and widely studied dark matter candidate is a weakly interacting massive particle (WIMP), which is a hypothetical stable particle, assumed to be a Majorana fermion with masses in the GeV-TeV range and weak-scale interactions Lee and Weinberg (1977); Jungman et al. (1996); Bertone et al. (2005). The most well motivated WIMP candidate is the lightest supersymmetric particle, which is usually the neutralino Ellis et al. (1984). Such a hypothetical particle has velocity averaged cross-section given by . This cross-section yields the correct relic abundance, and enables the WIMP to decouple from the rest of the universe at non-relativistic velocities, thereby providing all the right properties needed for a cold dark matter candidate Jungman et al. (1996); Bertone et al. (2005); Steigman et al. (2012). There are two kinds of experiments designed to look for signatures of WIMPs. The first type of search involves direct detection of WIMPs, where one looks for signatures of WIMP scattering with nuclei in an underground experiment Schumann (2019). The second search is referred to as indirect detection Cirelli (2012); Donato (2014), where one looks for signatures of WIMP annihilation into secondary particles such as such as neutrinos Desai et al. (2004), positrons Cholis et al. (2009), anti-protons Heisig (2021), radio waves Chan (2021), and gamma-rays Funk (2015).
@@
The Fermi Gamma-ray Space Telescope which was launched in 2008 is currently the most sensitive telescope for gamma-ray astronomy. The Large Area Telescope (LAT) is one of the two instruments onboard this detector. Fermi-LAT is sensitive to high energy gamma rays from various astrophysical sources. It is a pair-conversion telescope that is sensitive to photons between the energy range of 20 MeV to more than 300 GeV Atwood et al. (2009) and is sensitive to dark matter annihilation if the dark matter is a WIMP with masses in the GeV range, since the photon energy from the annihilation products should be within the detector sensitivity.
@@
A large number of searches have been done looking for dark matter annihilation in galaxy clusters using Fermi-LAT Ackermann et al. (2010); Ando and Nagai (2012); Huang et al. (2012); Han et al. (2012); Hektor et al. (2013); Anderson et al. (2016); Liang et al. (2016); Ackermann et al. (2015); Lisanti et al. (2018); Thorpe-Morgan et al. (2021); Shen et al. (2021); Di Mauro et al. (2023); Song et al. (2024). We briefly summarize these works. The very first search for dark matter annihilation from clusters was done using 11 months of Fermi-LAT data from six X-ray selected clusters in the HIFLUGCS catalog Ackermann et al. (2010). No significant gamma-ray detection was reported in this analysis. Subsequently, null results were reported based on a search from 49 galaxy clusters located at high galactic latitude using 2.8 years of Fermi-LAT data Ando and Nagai (2012). The most stringent upper limit from this analysis was from the Fornax cluster, given by Ando and Nagai (2012). Then, another search for dark matter annihilation and decay was done from eight nearby galaxy clusters (both individually and from a stacked analysis) using three years of Fermi-LAT data Huang et al. (2012). No evidence for gamma-ray emission at significance due to dark matter decay or annihilation was found from this analysis. Limits on the annihilation cross-section of around were then obtained Huang et al. (2012). A similar search for dark matter annihilation from three nearby clusters, namely, VIRGO, Coma, Fornax was done using 45 months of Fermi-LAT data Han et al. (2012). After removing the galactic and extragalactic gamma-ray foregrounds, a residual diffuse emission was found, which was attributed to additional point sources not included in the Fermi-LAT two-year catalog Han et al. (2012). After accounting for these, no significant extended emission was seen from the aforementioned three clusters and upper limits on the cross section for dark matter annihilation from the three clusters were obtained, which were more stringent than those obtained from Milky way dwarf galaxies Han et al. (2012). Subsequently, Hektor et al. (2013) found evidence for a double-peaked line at 110 and 130 GeV from 18 nearby brightest galaxy clusters with more than significance using 218 weeks of Fermi-LAT data, which they attributed to dark matter annihilation. However, this signal could not be confirmed in a subsequent follow-up analysis using enhanced exposure Anderson et al. (2016); Liang et al. (2016). A dedicated search for dark matter annihilation and point source emission from the VIRGO cluster was undertaken by the Fermi-LAT collaboration using three years of data, which reported null results Ackermann et al. (2015). This work then calculated limits on velocity averaged DM annihilation cross-section. Then in 2016, two independent groups did a search around the same time from the same set of 16 galaxy clusters from the HIFLUGCS sample with the largest -factors, using five and seven years of data respectively Anderson et al. (2016); Liang et al. (2016). The first analysis Anderson et al. (2016) did not find any significant excess in the energy range between 10 to 400 GeV and upper limits on the velocity averaged annihilation cross-section of around were obtained. The second analysis found a line at 43 GeV after stacking with a significance of Liang et al. (2016). However, since the backgrounds were not fully understood, they did not claim a detection and set a conservative limit on the the velocity averaged annihilation cross-section Liang et al. (2016). A search for dark matter annihilation from galaxy groups (lower mass analogs of galaxy clusters) having large -factors using 413 weeks of Fermi-LAT data Lisanti et al. (2018). No evidence for dark matter annihilations was found and thermal relic cross sections for dark matter masses below approximately 30 GeV to annihilation channel were excluded at 95% c.l. Then, another search for dark matter annihilation using 12 years of Fermi-LAT data from five nearby clusters (Centaurus, Coma, VIRGO, Fornax, and Perseus) reported null results and constraints on annihilation cross-section into three different channels were reported, which are given between , depending on the WIMP mass Thorpe-Morgan et al. (2021). A search for line and box-shaped signals was done in Shen et al. (2021) using 7.1 years of Fermi-LAT by stacking the same set of 16 clusters analyzed in Anderson et al. (2016); Liang et al. (2016). This work also found the 42 GeV line first reported in Liang et al. (2016), although its significance was reduced. Consequently , 95% c.l. upper limits on the thermally averaged DM annihilation cross-section to photons between were set Shen et al. (2021). About a year ago, Di Mauro et al. (2023) (D23, hereafter) used 12 years of Fermi-LAT data to look for dark matter annihilation and decay from 49 clusters selected from the HIFLUGCS sample. The stacking analysis revealed a signal at 2.5-3.0 significance. However, the best-fit values of the mass and cross-section are in tension with the null results obtained using dark matter searches from dwarf spheroidal galaxies. Therefore, it was concluded that the signal is most likely due to cosmic ray collisions with the gas and photon fields within the cluster Di Mauro et al. (2023). Most recently, a constraint on the lifetime of very heavy dark matter with masses between and GeV was set using 14 years of Fermi-LAT data, based on the analysis of seven clusters Song et al. (2024).
@@
Complementary to the aforementioned target searches from individual clusters or a stacked analysis, a cross-correlation analysis between the 9-year Fermi-LAT diffuse gamma-ray map and four different galaxy cluster catalogs was done to look for dark matter annihilations, which reported null results Tan et al. (2020). In additional to galaxy clusters, a large number of works have also looked for dark matter annihilations and decay from Galactic center, dwarf spheriodal galaxies as well as Milky way satellites.
@@
In this work, we look for WIMP dark matter annihilation from the galaxy clusters detected by the South Pole Telescope (SPT) 2500 sq. deg survey, which have been detected using the Sunyaev-Zeldovich (SZ) effect Sunyaev and Zeldovich (1972), as a followup to our previous work which looked for point-source gamma-ray emission using 15 years of Fermi-LAT data Manna and Desai (2024a). We mainly follow the same methodology as D23. In our analysis, we have assumed a flat CDM cosmology with and . This manuscript is structured as follows: In Section II, we define the cluster catalog utilized in our manuscript. Section III provides a comprehensive description of the halo and subhalo modelling. Subsequently, in Section IV, we detail the analysis process employed in our study. Finally, in Section V, we present our results and present our conclusions in Section VI.
@@ II SPT-SZ cluster catalog
For our analysis, we use the SZ selected cluster catalog detected by SPT, which provides a mass-limited sample. The SPT is a 10-meter telescope located at the South Pole that has imaged the sky at three different frequencies: 95 GHz, 150 GHz, and 220 GHz with an angular resolution of about 1 arcminute Carlstrom et al. (2011). SPT completed a 2500 square-degree survey between 2007 and 2011 to detect galaxy clusters using the SZ effect. This 2500 sq. degree SPT-SZ survey detected 677 confirmed galaxy clusters with SNR greater than 4.5, corresponding to a mass threshold of up to redshift of 1.8 Bleem et al. (2015a); Bocquet et al. (2019)111https://pole.uchicago.edu/public/data/sptsz-clusters/2500d_cluster_sample_Bocquet19.fits. The SPT cluster redshifts were obtained using a dedicated optical and infrared follow-up campaign from pointed imaging and spectroscopic observations Song et al. (2012); Ruel et al. (2014) in conjunction with data from optical imaging surveys such as the Blanco Cosmology Survey Desai et al. (2012) and Dark Energy Survey Saro et al. (2015). We have previously done a search for gamma-rays induced due to astrophysical processes from 300 clusters from this sample, ranked according to Manna and Desai (2024a). Here, denotes the total mass enclosed within a sphere with a mean density of 500 times the critical density of the universe at the cluster’s redshift, and represents the cluster’s redshift Bocquet et al. (2019). In this work, we conduct a search for gamma-ray emission from dark matter annihilation for a sample of 350 galaxy clusters from the aforementioned catalog, the clusters are ranked in decreasing order of their mass ratio, .
@@ III Dark Matter Analysis
where is the angle between the line-of-sight and the direction pointing toward the centre of the astrophysical system. In our study, the integration angle is considered as 0.1∘. In the context of modelling the dark matter distribution within galaxy clusters, a common approach is to consider the total dark matter density profile as the sum of two components Di Mauro et al. (2023):
@@ III.1 Modelling the Main Halo
The first step in obtaining the halo concentration involves choosing a mass proxy. For this purpose, we use , which is the cluster mass at a spherical over-density with a density 200 times the critical density of the Universe. We have obtained values from the SPT-SZ 2500 square-degree catalogue Bocquet et al. (2019); Bleem et al. (2015b). In accordance with Di Mauro et al. (2023); White (2001), we define the virial radius, which can be estimated from as follows:
@@
Next, we need to determine for every cluster. Although, an extensive weak lensing campaign has been undertaken for the SPT clusters for precise mass measurements and cosmological analyses Bocquet et al. (2019) (and references therein), no concentration measurements are available on a per cluster basis for the SPT sample. Therefore, we use the concentration-mass () relations from cosmological simulations for modelling the dark matter halo. Similar to D23, we use the concentration-mass relations from Sánchez-Conde and Prada (2014), which have been shown to match the observation results across a wide range of halo masses, from dwarf spheroidal galaxies to galaxy clusters. This mass-concentration relations allows us to obtain from the estimated for every SPT cluster.
@@ III.2 Modelling the subhalo
In the above equation, refers to the expected number of subhalos within the virial radius of main halo, and with is the probability distribution in each of the respective domains, normalized to 1. Here, corresponds to the volume of the main halo, pertains to the distribution of subhalo masses, and represents the subhalo concentration. Although, numerical cosmological simulations have significantly advanced our understanding of halo substructures Zavala and Frenk (2019), several questions still remain unresolved, such as the minimum mass for clump formation Colafrancesco et al. (2006), the effects of tidal stripping on subhalo survival Aguirre-Santaella et al. (2023), and the exact shape of subhalo DM density profiles Errani and Navarro (2021). These uncertainties impact the calculation of the DM-induced -ray flux. In this work, we follow the same prescription as D23 to account for each of the terms in Eq. 11, which we describe below:
Spatial distribution PDF (): Since the main halo is spherically symmetrical, the distribution of subhalos within it depends only on their distance from the central point of the host halo. It describes the probability of finding a subhalo at a given distance from the host halo centre and provides the spatial distribution of substructures in a host halo, which in our study are the galaxy clusters. We modelled it using the Einasto profile.

### v2

@@ Abstract
We search for dark matter annihilation from galaxy clusters in the energy range from 1-300 GeV using nearly 16 years of Fermi-LAT data. For this purpose, we use 350 galaxy clusters selected from the 2500 SPT-SZ survey. We model the dark matter distribution using the NFW profile for the main halo along with the Einasto profile for the substructure. The largest signal is seen for the cluster SPT-CL J2021-5257 with a significance of around . The best-fit dark matter mass and annihilation cross-section for this cluster are equal to GeV and for the annihilation channel. However, this central estimate is in conflict with the limits on annihilation cross-section from dwarf spheroidal galaxies, and hence cannot be attributed to dark matter annihilation. Three other clusters show significance between , whereas all the remaining clusters show null results. The most stringent 95% c.l. upper limit for the WIMP annihilation cross-section among all the clusters is from SPT-CL J0455-4159, viz. for GeV and annihilation channel.
@@ I Introduction
The dark matter problem is one of the most outstanding unsolved problems in Astrophysics Jungman et al. (1996); Bertone et al. (2005); Bertone and Hooper (2018); Bosma (2023). Although a large number of observations have pointed to a concordance model of the universe, which consists of about 25% cold dark matter Planck Collaboration et al. (2020), the identity of the cold dark matter candidate is unknown. One of the most well motivated and widely studied dark matter candidate is a weakly interacting massive particle (WIMP), which is a hypothetical stable particle, assumed to be a Majorana fermion with masses in the GeV-TeV range and weak-scale interactions Lee and Weinberg (1977); Jungman et al. (1996); Bertone et al. (2005). The most well motivated WIMP candidate is the lightest supersymmetric particle, which is usually the neutralino Ellis et al. (1984). Such a hypothetical particle has velocity averaged cross-section given by . This cross-section yields the correct relic abundance, and enables the WIMP to decouple from the rest of the universe at non-relativistic velocities, thereby providing all the right properties needed for a cold dark matter candidate Jungman et al. (1996); Bertone et al. (2005); Steigman et al. (2012). There are three kinds of experiments designed to look for signatures of WIMPs. The first type of search involves direct detection of WIMPs, where one looks for signatures of WIMP scattering with nuclei in an underground experiment Schumann (2019). The second search is referred to as indirect detection Zeldovich et al. (1980); Cirelli (2012); Donato (2014), where one looks for signatures of WIMP annihilation into secondary particles such as such as neutrinos Desai et al. (2004), positrons Cholis et al. (2009), anti-protons Heisig (2021), radio waves Chan (2021), and gamma-rays Funk (2015). Finally, one can also look for dark matter signatures at particle colliders such as LHC Boveia and Doglioni (2018).
@@
The Fermi Gamma-ray Space Telescope which was launched in 2008 is currently the most sensitive telescope for gamma-ray astronomy at GeV energies. The Large Area Telescope (LAT) is one of the two instruments onboard this detector. Fermi-LAT is sensitive to high energy gamma rays from various astrophysical sources. It is a pair-conversion telescope that is sensitive to photons between the energy range of 20 MeV to more than 300 GeV Atwood et al. (2009) and is sensitive to dark matter annihilation if the dark matter is a WIMP with masses in the GeV range, since the photon energy from the annihilation products should be within the detector sensitivity.
@@
A large number of searches have been done looking for dark matter annihilation in galaxy clusters using Fermi-LAT Ackermann et al. (2010); Ando and Nagai (2012); Huang et al. (2012); Han et al. (2012); Hektor et al. (2013); Anderson et al. (2016); Liang et al. (2016); Ackermann et al. (2015); Lisanti et al. (2018); Thorpe-Morgan et al. (2021); Shen et al. (2021); Di Mauro et al. (2023); Song et al. (2024). We briefly summarize these works. The very first search for dark matter annihilation from clusters was done using 11 months of Fermi-LAT data from six X-ray selected clusters in the HIFLUGCS catalog Ackermann et al. (2010). No significant gamma-ray detection was reported in this analysis. Subsequently, null results were reported based on a search from 49 galaxy clusters located at high galactic latitude using 2.8 years of Fermi-LAT data Ando and Nagai (2012). The most stringent upper limit from this analysis was from the Fornax cluster, given by for a WIMP mass of around 10 GeV Ando and Nagai (2012). Then, another search for dark matter annihilation and decay was done from eight nearby galaxy clusters (both individually and from a stacked analysis) using three years of Fermi-LAT data Huang et al. (2012). No evidence for gamma-ray emission at significance due to dark matter decay or annihilation was found from this analysis. Limits on the annihilation cross-section of around were then obtained Huang et al. (2012). A similar search for dark matter annihilation from three nearby clusters, namely, VIRGO, Coma, Fornax was done using 45 months of Fermi-LAT data Han et al. (2012). After removing the galactic and extragalactic gamma-ray foregrounds, a residual diffuse emission was found, which was attributed to additional point sources not included in the Fermi-LAT two-year catalog Han et al. (2012). After accounting for these, no significant extended emission was detected from the aforementioned three clusters, and upper limits on the cross section for dark matter annihilation were obtained, which were more stringent than those obtained from Milky way dwarf galaxies Han et al. (2012). Subsequently, Hektor et al. (2013) found evidence for a double-peaked line at 110 and 130 GeV from 18 nearby brightest galaxy clusters with more than significance using 218 weeks of Fermi-LAT data, which they attributed to dark matter annihilation. However, this signal could not be confirmed in a subsequent follow-up analysis using enhanced exposure Anderson et al. (2016); Liang et al. (2016). A dedicated search for dark matter annihilation and point source emission from the VIRGO cluster was undertaken by the Fermi-LAT collaboration using three years of data, which reported null results Ackermann et al. (2015). This work then calculated limits on velocity averaged DM annihilation cross-section. Then in 2016, two independent groups did a search around the same time from the same set of 16 galaxy clusters from the HIFLUGCS sample with the largest -factors, using five and seven years of data respectively Anderson et al. (2016); Liang et al. (2016). The first analysis did not find any significant excess in the energy range between 10 to 400 GeV and upper limits on the velocity averaged annihilation cross-section of around were obtained Anderson et al. (2016). The second analysis found a line at 43 GeV after stacking with a significance of Liang et al. (2016). However, since the backgrounds were not fully understood, they did not claim a detection and set a conservative limit on the the velocity averaged annihilation cross-section Liang et al. (2016). A search for dark matter annihilation from galaxy groups (lower mass analogs of galaxy clusters) having large -factors was carried out using 413 weeks of Fermi-LAT data Lisanti et al. (2018). No evidence for dark matter annihilations was found and thermal relic cross sections for dark matter masses below approximately 30 GeV to annihilation channel were excluded at 95% c.l. Then, another search for dark matter annihilation using 12 years of Fermi-LAT data from five nearby clusters (Centaurus, Coma, VIRGO, Fornax, and Perseus) reported null results and constraints on annihilation cross-section into three different channels were reported, which are between , depending on the WIMP mass Thorpe-Morgan et al. (2021). A search for line and box-shaped signals was done in Shen et al. (2021) using 7.1 years of Fermi-LAT data by stacking the same set of 16 clusters analyzed in Anderson et al. (2016); Liang et al. (2016). This work also found the 42 GeV line first reported in Liang et al. (2016), although its significance was reduced. Consequently, 95% c.l. upper limits on the thermally averaged DM annihilation cross-section to photons between were set Shen et al. (2021). About a year ago, Di Mauro et al. (2023) (D23, hereafter) used 12 years of Fermi-LAT data to look for dark matter annihilation and decay from 49 clusters selected from the HIFLUGCS sample. The stacking analysis revealed a signal at 2.5-3.0 significance. However, the best-fit values of the mass and cross-section are in tension with the null results obtained using dark matter searches from dwarf spheroidal galaxies. Therefore, it was concluded that the signal is most likely due to cosmic ray collisions with the gas and photon fields within the cluster Di Mauro et al. (2023). Most recently, a constraint on the lifetime of very heavy dark matter with masses between and GeV was set using 14 years of Fermi-LAT data, based on the analysis of seven clusters Song et al. (2024).
@@
Complementary to the aforementioned target searches from individual clusters or a stacked analysis, a cross-correlation analysis between the 9-year Fermi-LAT diffuse gamma-ray map and four different galaxy cluster catalogs was done to look for dark matter annihilations, which reported null results Tan et al. (2020). In addition to galaxy clusters, a large number of works have also looked for dark matter annihilations and decay from Galactic center, dwarf spheriodal galaxies as well as Milky way satellites Murgia (2020).
@@
In this work, we look for WIMP dark matter annihilation from galaxy clusters detected by the South Pole Telescope (SPT) 2500 square deg survey, which have been detected using the Sunyaev-Zeldovich (SZ) effect Sunyaev and Zeldovich (1972), as a follow-up to our previous work which looked for point-source gamma-ray emission using 15 years of Fermi-LAT data Manna and Desai (2024a). We mainly follow the same methodology as D23. In our analysis, we have assumed a flat CDM cosmology with and . This manuscript is structured as follows. In Section II, we define the cluster catalog utilized in our manuscript. Section III provides a comprehensive description of the halo and subhalo modelling. Subsequently, in Section IV, we detail the analysis process employed in our study. Finally, in Section V, we present our results. The upper limits for all clusters can be found in Sect. VI. We present our conclusions in Section VII.
@@ II SPT-SZ cluster catalog
For our analysis, we use the SZ selected cluster catalog detected by SPT, which provides a mass-limited sample. The SPT is a 10-meter telescope located at the South Pole that has imaged the sky at three different frequencies: 95 GHz, 150 GHz, and 220 GHz with an angular resolution of about 1 arcminute Carlstrom et al. (2011). SPT completed a 2500 square degree survey between 2007 and 2011 to detect galaxy clusters using the SZ effect. This survey detected 677 galaxy clusters with SNR greater than 4.5, corresponding to a mass threshold of up to redshift of 1.8 Bleem et al. (2015a); Bocquet et al. (2019)111https://pole.uchicago.edu/public/data/sptsz-clusters/2500d_cluster_sample_Bocquet19.fits. The SPT cluster redshifts were obtained using a dedicated optical and infrared follow-up campaign from pointed imaging and spectroscopic observations Song et al. (2012); Ruel et al. (2014) in conjunction with data from optical imaging surveys such as the Blanco Cosmology Survey Desai et al. (2012) and Dark Energy Survey Saro et al. (2015). We have previously done a search for gamma-rays induced due to astrophysical processes from 300 clusters from this sample, ranked according to Manna and Desai (2024a). Here, denotes the total mass enclosed within a sphere with a mean density of 500 times the critical density of the universe at the cluster’s redshift, and represents the cluster’s redshift Bocquet et al. (2019). In this work, we conduct a search for gamma-ray emission from dark matter annihilation for a sample of 350 galaxy clusters from the aforementioned catalog, after ranking them in decreasing order of their mass ratio, .
@@ III Dark Matter Analysis
where is the angle between the line-of-sight and the direction pointing toward the centre of the cluster. We have chosen to be 0.2∘ similar to size of the spatial bins used for the analysis. In the context of modelling the dark matter distribution within galaxy clusters, a common approach is to consider the total dark matter density profile as the sum of two components Di Mauro et al. (2023):
@@ III.1 Modelling the Main Halo
The first step in obtaining the halo concentration involves choosing a mass proxy. For this purpose, we use , which is the cluster mass at a spherical overdensity, which is about 200 times the critical density of the Universe. We have obtained values from the SPT-SZ 2500 square-degree catalogue Bocquet et al. (2019); Bleem et al. (2015b). In accordance with Di Mauro et al. (2023); White (2001), we define the virial radius, which can be estimated from as follows:
@@
Next, we need to determine for every cluster. Although, an extensive weak lensing campaign has been undertaken for SPT clusters for precise mass measurements and cosmological analyses Bocquet et al. (2019) (and references therein), no concentration measurements are available on a per cluster basis. Therefore, we use the concentration-mass () relations from cosmological simulations for modelling the dark matter halo. Similar to D23, we use the concentration-mass relations from Sánchez-Conde and Prada (2014), which have been shown to match the observational results across a wide range of halo masses from dwarf spheroidal galaxies to galaxy clusters. This mass-concentration relation allows us to obtain from the estimated for every SPT cluster.
@@ III.2 Modelling the subhalo
In the above equation, refers to the expected number of subhalos within the virial radius of main halo, and with is the probability distribution in each of the respective domains, normalized to 1. Here, corresponds to the volume of the main halo, pertains to the distribution of subhalo masses, and represents the subhalo concentration. Although, numerical cosmological simulations have significantly advanced our understanding of halo substructures Zavala and Frenk (2019), several questions still remain unresolved, such as the minimum mass for clump formation Colafrancesco et al. (2006), the effects of tidal stripping on subhalo survival Aguirre-Santaella et al. (2023), and the exact shape of the subhalo DM density profiles Errani and Navarro (2021). These uncertainties impact the calculation of the DM-induced -ray flux. In this work, we follow the same prescription as D23 to account for each of the terms in Eq. 11, which we describe below:
Spatial distribution PDF (): Since the main halo is spherically symmetric, the distribution of subhalos within it depends only on their distance from the central point of the host halo. It describes the probability of finding a subhalo at a given distance from the host halo centre and provides the spatial distribution of substructures in a host halo, which in our study are the galaxy clusters. We modelled it using the Einasto profile.